OpenStudy (anonymous):

Solve: T(3) = 3/√3^2+9 the square root covers the 3 squared plus nine. Thank you!

OpenStudy (anonymous):

wait, is there meant to be an t, x or some kind of pronumeral in there?

OpenStudy (anonymous):

it was T(L)=3/√L^2+9, T(3)= Originally

OpenStudy (anonymous):

\[T(L)=3\div \sqrt{L^2+9}\]

OpenStudy (anonymous):

yes kk. T(3) = \[3 \div \sqrt{L^2 + 9}\] = 3 / sqrt 3^2 + 9 = 3 / sqrt 18 = 3/ 3sqrt 2. = 1/sqrt 2

OpenStudy (anonymous):

sorry, I wish that the equation symbol and typing all the symbols would have been more smooth and easy to do, but in fact it's not.

OpenStudy (anonymous):

yeah i know. my teacher said the answer is \[\sqrt{2}\div2\]

OpenStudy (mertsj):

\[T(3)=\frac{3}{\sqrt{3^2+9}}=\frac{3}{\sqrt{18}}=\frac{3}{3\sqrt{2}}=\frac{1}{\sqrt{2}}=\frac{\sqrt{2}}{2}\]

OpenStudy (anonymous):

How did you go from 1/√2 to √2/2?

OpenStudy (mertsj):

\[\frac{1}{\sqrt{2}}\times\frac{\sqrt{2}}{\sqrt{2}}=\frac{\sqrt{2}}{2}\]